KyledaGreat:

Consider the following relation. x=2|y|−1 Find four points contained in the inverse. Express your values as an integer or simplified fraction. {( , ), ( , ), ( , ), ( , )}''.

surjithayer:

2|y|=x+1 flip x and y 2|x|=y+1 y=2|x|-1 x=-1,y=2|-1|-1=2-1=1 (-1,1) x=1,y=2|1|-1=2-1=1 (1,1) x=-2,y=2|-2|-1=2(2)-1=4-1=3 (-2,3) x=2,y=2|2|-1=4-1=3 (2,3) points are (-1,1),(1,1),(-2,3),(2,3)